If you are traveling 55 miles per hour and you need to go 44 miles how many
minuets will it take to travel the 44 miles?

Answer by mananth(16949)

Distance = 44 miles
speed = 55 mph
t=d/r
t=44/55
t=4/5 hours
4/5 * 60 = 48 minutes